Lorrie Morgan places a 29 ft ladder against the side of a building with the bottom of the ladder 20 ft away from the building How high up on the wall does the ladder reach?

Answer by wilft1(217)

a^2+b^2=c^2
a^2+400=841
-400 -400
a^2=441
a=21
THE LADDER GOES 21 FEET UP THE WALL